a car with a mass of 900 kg accelerates along the road at 3m/s2. a frictional force of 600 n is acting on the wheels. what force is produced by the car's engine?


1
Expert's answer
2021-03-23T08:14:17-0400

Let's first find the net force of the car from the Newton's Second Law of Motion:


"F_{net}=ma=900\ kg\cdot3\ \dfrac{m}{s^2}=2700\ N."

Then, we can find the force produced by the car's engine:


"F_{engine}=F_{net}-F_{fr}=2700\ N-600\ N=2100\ N."